write a step-by-step reaction mechanism for collagen cross-linking via curbodiimide route?

Collagen Catoomide R2 CRI NH Collag Collagen-CO,H NH, Collagen Ro
Answer #1

Mechanism Но-N Collagen-cooH + N-Collagen + R-N=c=N-R2 Collagen--NHollogen Mechanism Collagen Collagen-E-07HK R-N=c=NR2 RCNR